45) Find the integral surface of the equation x(y ^ 2 + z) * p - y(x ^ 2 + z) * q = (x ^ 2 - y ^ 2) * z whic the straight line x + y = 0, z = 1


1
Expert's answer
2022-04-05T13:03:13-0400

"x(y ^ 2 + z)\frac{\partial z}{dx}- y(x ^ 2 + z)\frac{\partial z}{dy}=(x ^ 2 - y ^ 2) z"

"x + y = 0, z = 1"

The auxiliary equations is:

"\frac{dx}{x(y ^ 2 + z)}=\frac{dy}{- y(x ^ 2 + z)}=\frac{dz}{(x ^ 2 - y ^ 2) z}"

A first characteristic equation comes from

"\frac{xdx+ydy}{x^2(y ^ 2 + z)- y^2(x ^ 2 + z)}=\frac{dz}{(x ^ 2 - y ^ 2) z}"

"\frac{xdx+ydy}{(x ^ 2 - y ^ 2) z}=\frac{dz}{(x ^ 2 - y ^ 2) z}"

"xdx+ydy=dz"

"\frac{x^2+y^2}{2}=z+\frac{C_1}{2}"

"C_1=x^2+y^2-2z"

A second characteristic equation comes from

"\frac{ydx+xdy}{xy(y ^ 2 + z)- xy(x ^ 2 + z)}=\frac{dz}{(x ^ 2 - y ^ 2) z}"

"\frac{ydx+xdy}{-xy(x ^ 2 - y ^ 2) }=\frac{dz}{(x ^ 2 - y ^ 2) z}"

"\frac{d(xy)}{-xy}=\frac{dz}{ z}"

"-\ln{|xy|}+\ln{C_2}=\ln {|z|}"

"\frac{C_2}{xy}=z"

"C_2=xyz"

General solution of the PDE on the form of implicit equation:

"\Phi(C_1,C_2)=0".

"C_1=x^2+y^2-2z=x^2+y^2+2xy-2xy-2z=""(x+y)^2-2xy-2z" (1)

"C_2=xyz" (2)

Substitute (2) into (1):

"C_1=(x+y)^2-2\frac{C_2}{z}-2z"

Use "x + y = 0, z = 1" to exclude "C_1" and "C_2" :

"C_1=(0)^2-2\frac{C_2}{1}-2\cdot1=-2C_2-2"

"-2C_2-2=x^2+y^2-2z"

"-2xyz-2=x^2+y^2-2z"

"z=\frac{x^2+y^2}{2}+xyz+1"

Answer: "z=\frac{x^2+y^2}{2}+xyz+1" .